# Andreas Petrou-Zeniou

### v3

@@ 6 Application: Neighborhood Effects Revisited
In this section, we revisit the studies of Chetty et al. (2018) and Bergman et al. (2024) on the geographic nature of economic mobility. These studies have garnered substantial interest in the selective inference and multiple testing literatures, notably in Andrews et al. (2023) and Mogstad et al. (2023). Chetty et al. (2018) construct the so-called opportunity atlas, a dataset containing correlational estimates of the effect of childhood neighborhood in adulthood. In particular, Chetty et al. (2018) provide a measure of economic mobility by reporting a child’s expected earnings in adulthood (as a percentile) conditional on growing up in a given census tract and on a given parental income percentile. Chetty et al. (2018) also report analogous economic mobility estimates at the commuting zone (CZ) level. In a study motivated by the findings of Chetty et al. (2018), Bergman et al. (2024) study the effects of an informational intervention on the residential decisions of low-income housing voucher recipients in Seattle via the Creating Moves to Opportunity (CMTO) program. In particular, Bergman et al. (2024) use the economic mobility estimates of Chetty et al. (2018) to identify a set of high-opportunity tracts which they advertised to treated households in the CMTO program. We seek to study, in a set of exercises related to Andrews et al. (2022) and Mogstad et al. (2023), whether the CMTO can be expected to provide positive, long-run effects on the earnings of children in the treatment group, and relatedly whether the selection of high opportunity neighborhoods reflects noise as opposed to signal.
@@
Our analyses are closely related to a larger literature in theoretical econometrics studying the mobility estimates of Chetty et al. (2018) and Bergman et al. (2024). Andrews et al. (2023) study the CMTO program of Bergman et al. (2024), showing that there exists a statistically significant, positive difference in the average mobility of high-opportunity tracts and the mobility of the average housing voucher recipient. Mogstad et al. (2023) show that one can say little about the relative ranks of census tracts or commuting zones according to economic mobility. Our analysis imposes a more strict coverage criterion than those of Andrews et al. (2023), but a less strict coverage criterion than that of Mogstad et al. (2023).
@@ 6.1 Empirical Findings
The CMTO program has attracted substantial interest in the selective inference literature within econometrics. We seek to provide further insights on tract-level effects, building on the closely related work of Mogstad et al. (2023). In particular, Mogstad et al. (2023) study the problem of ranking tracts according to their true economic mobility effects. In a particularly stark finding, Mogstad et al. (2023) find that one cannot reject the possibility that the bottom-ranked tract in Seattle, according to estimated economic mobility, lies in the top third of tracts according to true economic mobility. As a result, Mogstad et al. (2023) conclude that:
@@
“This method is conservative because it assumes that the econometrician is comparing all tracts in LA county (whereas in practice we focused on Watts given its well-known history of poverty and violence) and because it controls the family wise error rate (i.e., it requires that the probability that one or more of the millions of pairwise comparisons is wrong is less than 5%).”
In our analysis of tract-level effects, we address the former point. In particular, we seek to study tract-level effects for certain tracts of interest alone, namely high-opportunity tracts. Our methods are well-suited to problems where focusing power on units of interest is desirable. Our two-step approach to inference focuses power onto selected tracts, while correcting for the winners’ curse induced by selection. Our simulations demonstrate that our approach reduces over-coverage substantially relative to projection and provides substantially shorter confidence regions. However, we find that even when focusing inference on tracts of interest, the evidence on tract-level effects remains murky. Some aggregation of effects across tracts or loosening of the simultaneous coverage requirement may be necessary for informative inference.
In the CMTO program, Bergman et al. (2024) designate neighborhoods as high-opportunity according to their ranks. In particular, they label the top 20% and top 40% of urban and non-urban tracts, respectively, in the Seattle commuting zone as high-opportunity. This roughly corresponds to the top third of tracts in the commuting zone.
In our first analysis of the CMTO program, we compare the economic mobility estimates at each selected, high-opportunity tract with an estimate of the average economic mobility of housing voucher recipients. To be precise, we allow $X_{j}$ to be the economic mobility estimate for tract $j$ in the set of all census tracts $J$ in the Seattle commuting zone, or alternatively the set of all census tracts in urban Seattle. For each tract $j$, we also observe the number of housing voucher recipients residing in $j$, which we denote by $c_{j}$. We define $Y_{j}$ as follows:
@@
$Y_{j}$ can be thought of as a point estimate of the gain from moving to tract $j$ for the average housing voucher recipient.
In this first exercise, we take $J$ to be the set of all urban Seattle tracts. We let $p=132$ be the number of tracts in urban Seattle, and we take $R$ to be $\{1,\ldots,\lfloor p/5\rfloor=26\}$. We are concerned with inference for the means $\mu_{Y,\jmath}$ for $\jmath$ in $\hat{J}_{R}$. In the figure below, we plot lower and upper confidence bounds for the mobility gains in selected Seattle tracts. We also apply the same exercise to Cleveland to demonstrate heterogeneity in findings between urban areas.
@@
In urban Seattle, we fail to reject the possibility of null effects at 92% of selected tracts. In urban Cleveland, we fail to reject the possibility of null effects in a comparatively small 35% of selected tracts. Our two-step confidence regions are essentially equivalent to projection in urban Seattle (providing being 1.2% shorter confidence intervals on tract-level effects), while in urban Cleveland, our two-step confidence regions are a modest but non-negligible 3.5% shorter than the projection confidence regions.
@@
$\bar{\mu}_{Y,\hat{J}_{R}}:=\frac{1}{\lfloor p/5\rfloor}\sum_{\jmath\in\hat{J}_{R}}\mu_{Y,\jmath}$
Their confidence region for $\bar{\mu}_{Y,\hat{J}_{R}}$ lies above zero, implying that they can reject the possibility of a null effect on economic mobility on aggregate. Our analysis differs from that of Andrews et al. (2023) in that we are concerned with constructing a confidence region satisfying simultaneous coverage of the $\mu_{Y,\jmath}$. Nevertheless, we believe that our analysis is complementary to that of Andrews et al. (2022), in that it provides insight into which of the selected tracts most credibly drive the positive aggregate effects. In contrast to Andrews et al. (2022), we find that we cannot reject the possibility of null neighborhood effects at most of the selected census tracts. Given both the results in Mogstad et al. (2023), and the simple fact that the variance of a sample mean is less than the variance of the constituent random variable forming the mean, these results are not surprising.
Our results suggest that, while one can draw informative conclusions on the aggregate effects of the CMTO program on economic mobility, these results are less clear when disaggregating. Moreover, while it is certainly plausible and in fact likely that treated households in the CMTO program move to a high-opportunity tract uniformly at random or approximately so, our tract-level confidence regions allow for inference robust to violations of such movement patterns. [^5] Of course, in such an analysis, we assume that conditional on moving to a given high-opportunity tract, a treated household selects an individual house or apartment within that tract uniformly at random. After all, one can imagine further disaggregation of the neighborhood mobility estimates of Chetty et al. (2018). While our analysis is not robust to such concerns, we address the first-order problem of non-uniform movement patterns between tracts. Our findings suggest that, in a worst-case, we cannot reject the possibility of the CMTO program having zero or even negative effects on treated families’ economic mobility outcomes. [^6] Our failure to reject could be attributed to the conservativeness of our coverage criterion. As Chetty et al. (2018), such familywise error rate (FWER) controlling procedures tend to be quite conservative. We may, for example, opt for control of the false coverage rate as in Benjamini and Yekutieli (2005). Developing selective inference methods that satisfy a false coverage rate condition, rather than a simultaneous coverage condition, is an interesting area for future research.
@@
Figure 3: Proportion of selected tracts for which we fail to reject the possibility of a null effect on economic mobility relative to the average housing voucher recipient, by commuting zone. We provide results for all top-50 CZs by population, using both the two-step and projection confidence sets.
In some commuting zones, our two-step confidence regions manage to reject the possibility of null effects for a larger proportion of selected top-third tracts than projection. Moreover, in many commuting zones, we are able to reject the possibility of null effects in over half of the selected tracts. These results suggest that our approach to inference is reasonably powerful, and can provide informative inference relative to projection.
In our final exercise on tract-level effects, we study pairwise comparison of low and high-opportunity tracts. In particular, we consider the thought experiment of a household moving from an arbitrary low-opportunity to an arbitrary high-opportunity tract, and seek to study how often we can reject the possibility of a null effect on economic mobility associated with this move. As discussed in 5.4, this problem is closely related to the problem of inference for ranks. However, our analysis is notably different from that of Mogstad et al. (2023), since we restrict attention to pairwise comparisons of the tracts of interest, rather than to all pairwise comparisons. Consequently, in the analyses that follow, we can conduct inference on the effects of moving from low to high opportunity tracts, without enforcing coverage on the effects of moving between tracts that are not of interest.
@@
CZ Top-$k$ Bottom-$k$ Fail to Reject Lowest LCB Highest LCB TS % Shorter than
Two-Step Two-Step Two-Step Projection
Seattle 26 26 87% [-0.33, 0.59] [0.077, 0.50] 0%
26 1 100% [-0.19, 0.60] [-0.055, 0.66] 0.3%
1 1 100% [-0.055, 0.66] [-0.055, 0.66] 0.3%
Cleveland 34 34 31% [-0.076, 0.59] [0.11, 0.49] 0%
34 1 9% [-0.056, 0.38] [0.11, 0.49] 0%
1 1 0% [0.12, 0.48] [0.12, 0.48] 3.4%
@@
In the table below, we replicate our analysis of the mobility effects of moving from low to high opportunity areas at the commuting zone level. Our results on mobility effects at the commuting zone level are more conclusive than our findings at the tract-zone level. This is in part due to the simple fact that the standard errors on CZ level effects are smaller than those on tract level effects. We find that, for the majority of high and low-opportunity commuting zone pairs, we can indeed conclude that there exists a difference in the true mobilities of both CZs. Moreover, in this application our two-step confidence regions are modestly shorter than projection. However, out two-step confidence regions allow us to make substantially more rejections. In particular, projection fails to reject the no-difference null between 1% and 26% more often than two-step.
@@
Top Bottom % Fail to Reject Lowest LCB Highest LCB TS % Shorter than % Fewer # Additional
Two-Step Two-Step Two-Step Projection Fail to Reject Rejections
50% 50% 18.5% [-0.28, 0.35] [0.33, 0.44] 0.7% 1.0% 267
33% 33% 8.2% [-0.25, 0.50] [0.33, 0.44] 0.9% 1.8% 92
20% 20% 4.7% [-0.19, 0.52] [0.34, 0.44] 4.4% 12.2% 140
10% 10% 3.5% [-0.14, 0.55] [0.34, 0.44] 7.9% 35.6% 103
33% 50% 12.0% [0.28, 0.35] [0.33, 0.44] 0.8% 1.3% 149
20% 50% 8.3% [-0.22, 0.38] [0.33, 0.44] 2.1 % 4.2% 198
10% 50% 7.1% [-0.18, 0.41] [0.34, 0.44] 5.0% 7.7% 159
33% 67% 19.9% [-0.28, 0.35] [0.33, 0.44] 0.7% 1.0% 247
20% 80% 19.1% [-0.25, 0.29] [0.33, 0.44] 1.9% 2.9% 500
10% 90% 17.6% [-0.22, 0.31] [0.34, 0.44] 4.7% 6.9% 633
@@
Our findings suggest a more complicated story. We are able to distinguish high and low-opportunity commuting zones with high frequency. The methods of Mogstad et al. (2023) depend on confidence regions for all pairwise comparisons. On the other hand, our methods focus only on pairwise comparisons of high and low-medium mobility commuting zones. As our comparison with the analogous projection approaches demonstrates, this difference translates to real gains in power. Moreover, Mogstad et al. (2023) use pairwise comparisons to construct confidence regions for ranks, likely losing power in the process. Finally, we do not construct confidence regions for ranks based on the confidence regions for selected, pairwise differences summarized above, providing a further explanation for the differences in our conclusions.
@@ 7 Simulation Study
In this section, we conduct an extended simulation study comparing the projection and two-step approaches to inference on multiple winners across a broad range of simulation designs. We demonstrate that the two-step approach to inference on multiple winners outperforms the projection approach almost uniformly. This outperformance is most pronounced in instances where the set of asymptotic winners is a proper subset of the set of all possible selections. We observe improvements in absolute coverage probability of between three and 91%. We also observe improvements in mean confidence set volume of up to 34%.
@@
[…]
@@
1 10 100 1000 10000
A Projection 0.978 0.989 0.991 0.991 0.992
Two-Step 0.965 0.972 0.966 0.955 0.955
Zoom 0.979 0.989 0.977 0.951 0.950
Locally Simultaneous 0.978 0.980 0.971 0.955 0.955
B Projection 0.984 0.989 0.993 0.996 0.995
Two-Step 0.968 0.965 0.969 0.970 0.960
Zoom 0.985 0.989 0.985 0.985 0.971
Locally Simultaneous 0.980 0.977 0.979 0.976 0.962
C Projection 0.976 0.977 0.976 0.976 0.976
Two-Step 0.962 0.962 0.962 0.962 0.962
Zoom 0.977 0.977 0.977 0.977 0.977
Locally Simultaneous 0.979 0.979 0.979 0.979 0.979
D Projection 0.975 0.988 0.991 0.991 0.991
Two-Step 0.964 0.976 0.970 0.970 0.969
Zoom 0.976 0.988 0.979 0.976 0.976
Locally Simultaneous 0.977 0.983 0.978 0.978 0.979
@@
We find that, between the models described above, the two-step approach substantially outperforms the projection approach when the set of winners $J_{R}(P)$ is clear, and specifically a proper subset of $J$ such that $J\setminus J_{R}(P)$ is large. Moreover, in intermediate cases when the set of winners is moderately clear, the two-step approach outperforms the approach of Zrnic and Fithian (2024a), which is based on the zoom test. However, the approach of Zrnic and Fithian (2024a) more closely matches the desired 95% coverage level in the case of the clear winner than our two step approach. In general, the two step approach also outperforms the approach of Zrnic and Fithian (2024b). Quantitatively, we have that in the four simulations presented above, the two-step approach to inference can reduce absolute overcoverage error by up to 88% relative to projection inference, 50% relative to locally simultaneous inference, and 56% relative to the zoom test. Across all simulations, including those in appendix 13, the two-step approach to inference reduces overcoverage error by up to 96% relative to projection inference.
@@
1 10 100 1000 10000
A Two-Step 0.937 0.889 0.835 0.780 0.780
Zoom 1.003 0.966 0.872 0.786 0.763
Locally Simultaneous 1.013 0.950 0.879 0.781 0.780
B Two-Step 0.922 0.863 0.824 0.784 0.730
Zoom 1.003 0.962 0.900 0.842 0.758
Locally Simultaneous 1.000 0.926 0.888 0.827 0.731
C Two-Step 0.937 0.937 0.937 0.937 0.937
Zoom 1.003 1.003 1.003 1.003 1.003
Locally Simultaneous 1.014 1.014 1.014 1.014 1.014
D Two-Step 0.938 0.928 0.846 0.846 0.846
Zoom 1.003 0.995 0.914 0.873 0.873
Locally Simultaneous 1.014 1.003 0.887 0.886 0.886
@@
Moreover, the table above demonstrates that, over a wide range of data generating processes, the two-step approach to inference provides tighter confidence regions than the projection, zoom, and locally simultaneous approaches to inference. Indeed, the interval lengths of the two-step approach to inference may be up to 27% shorter than projection inference. Moreover, two-step inference may be up to 11% shorter than inversions of the zoom test and up to 8% shorter than two-step inference.
@@
In general, we recommend choosing $\beta=\alpha/10$. As our simulation results demonstrate, the two-step approach to inference performs quite well under such a choice of $\beta$. This choice is based on certain heuristics on the role of $\beta$ in two step inference. Choosing $\beta=0$ renders the two-step approach equivalent to the projection approach. The parameter $\beta$ allows us to model the errors of the maximum, rather than all errors simultaneously. A choice of $\beta$ close to zero limits our ability to restrict inference in the manner described above, but also requires a very minimal Bonferroni correction. A choice of $\beta$ close to $\alpha$ provides a tighter restriction on the indices of interest, at the expense of a larger Bonferroni correction. Our suggested choice of $\beta$ is the same as in Andrews et al. (2023) and Romano et al. (2014).
8 Conclusion
In this paper, we consider a generalization of the inference on winners setting described in Andrews et al. (2023), which we dub the inference on multiple winners setting. We review the existing approaches to inference on winners, and demonstrate that they are either not appropriate in the inference on multiple winners setting, or overly conservative. We therefore propose a novel two-step approach to inference on multiple winners based on the two-step approach to testing moment inequalities from Romano et al. (2014). We demonstrate both theoretically and via simulation that this two-step approach asymptotically outperforms existing approaches for a broad range of data generating processes where the winner is asymptotically clear.
We consider several empirical settings where the inference on multiple winners setting may be relevant. We revisit the JOBSTART demonstration of Cave et al. (1993) and Miller et al. (2005). We consider the thought experiment where retraining sites are selected for replication based on a statistical significance cutoff rule, meaning that multiple sites are selected with frequency ranging from 17% to 52%. Applying our two-step approach to inference in this setting provides confidence sets that are 8% shorter than projection approaches. We also revisit the Creating Moves to Opportunity program from Bergman et al. (2024), finding that we cannot reject the possibility of null neighborhood effects in many of the program’s selected census tracts, though this finding varies with the commuting zone studied. In this setting, we find that the two-step confidence set may provide tighter intervals on neighborhood effects, relative to the projection confidence set. Finally, we demonstrate via simulation that two-step approaches can outperform projection approaches by as much as 96% in overcoverage error and 27% in confidence set length. Thanks to these substantial improvements on key confidence set properties, derived both analytically and via simulation, we recommend using the two-step approach in settings where selection occurs on multiple dimensions, and a hybridized two-step approach where selection is one-dimensional, as in Andrews et al. (2023).

### v4

@@ 6 Application: Neighborhood Effects Revisited
In this section, we revisit the studies of Chetty et al. (2025) and Bergman et al. (2024) on the geographic nature of economic mobility. These studies have garnered substantial interest in the selective inference and multiple testing literatures, notably in Andrews et al. (2023) and Mogstad et al. (2023). Chetty et al. (2025) construct the Opportunity Atlas, a dataset containing correlational estimates of the effect of childhood neighborhood in adulthood. In particular, Chetty et al. (2025) provide a measure of economic mobility by reporting a child’s expected earnings in adulthood (as a percentile) conditional on growing up in a given census tract with parents whose earnings fall in a particular income percentile. Chetty et al. (2025) also report analogous economic mobility estimates at the commuting zone (CZ) level. In a study motivated by the findings of Chetty et al. (2025), Bergman et al. (2024) report the effects of an informational intervention on the residential decisions of low-income housing voucher recipients in Seattle via the Creating Moves to Opportunity (CMTO) program. In particular, Bergman et al. (2024) use the economic mobility estimates of Chetty et al. (2025) to identify a set of high-opportunity tracts which they advertised to treated households in the CMTO program. We seek to study, in a set of exercises related to Andrews et al. (2022) and Mogstad et al. (2023), whether the CMTO can be expected to provide positive, long-run effects on the earnings of children in the treatment group, and relatedly whether the selection of high opportunity neighborhoods reflects noise as opposed to signal.
@@
Our analyses are closely related to a larger literature in selective inference and multiple testing studying the mobility estimates of Chetty et al. (2025) and Bergman et al. (2024). Andrews et al. (2023) study the CMTO program of Bergman et al. (2024), showing that there exists a statistically significant, positive difference in the average mobility of high-opportunity tracts and the mobility of the average housing voucher recipient. Mogstad et al. (2023) show that one can say little about the relative ranks of census tracts or commuting zones according to economic mobility. Our analysis imposes a more strict coverage criterion than those of Andrews et al. (2023), but a less strict coverage criterion than that of Mogstad et al. (2023).
@@ 6.1 Empirical Findings
The CMTO program has attracted substantial interest in the selective inference and multiple testing literatures. We seek to provide further insights on tract-level effects, building on the work of Mogstad et al. (2023). In particular, Mogstad et al. (2023) study the problem of ranking tracts according to their true economic mobility effects. In a particularly stark finding, Mogstad et al. (2023) find that one cannot reject the possibility that the bottom-ranked tract in Seattle, according to estimated economic mobility, lies in the top third of tracts according to true economic mobility. As a result, Mogstad et al. (2023) conclude that:
@@
“This method is conservative because it assumes that the analyst is comparing all tracts in LA county (whereas in practice we focused on Watts given its well-known history of poverty and violence) and because it controls the family wise error rate (i.e., it requires that the probability that one or more of the millions of pairwise comparisons is wrong is less than 5%).”
In our analysis of tract-level effects, we focus power on selected tracts to address the former point. In particular, we seek to study tract-level effects for certain tracts of interest alone, namely high-opportunity tracts. Our two-step approach to inference focuses power onto selected tracts, while correcting for the winners’ curse induced by selection. Our simulations demonstrate that our approach reduces over-coverage substantially relative to projection and provides substantially shorter confidence regions. However, we find that even when focusing inference on tracts of interest, the evidence on tract-level effects remains murky. Some aggregation of effects across tracts or loosening of the simultaneous coverage requirement may be necessary for informative inference.
In the CMTO program, Bergman et al. (2024) designate a subset of Seattle neighborhoods as high-opportunity according to their ranks. In particular, they label the top 20% and top 40% of urban and non-urban tracts, respectively, in the Seattle commuting zone as high-opportunity. This roughly corresponds to the top third of tracts in the commuting zone.
In our first analysis of the CMTO program, we compare the economic mobility estimates in each selected high-opportunity tract with an estimate of the average economic mobility of housing voucher recipients. To be precise, we allow $X_{j}$ to be the economic mobility estimate for tract $j$ in the set of all census tracts $J$ in the Seattle commuting zone, or alternatively the set of all census tracts in urban Seattle. For each tract $j$, we also observe the number of housing voucher recipients residing in $j$, which we denote by $c_{j}$. We define $Y_{j}$ as follows:
@@
$Y_{j}$ is an estimate of the expected gain or loss from moving the average housing voucher recipient to tract $j$.
In this first exercise, we take $J$ to be the set of all urban Seattle tracts. We let $p=132$ be the number of tracts in urban Seattle, and we take $R$ to be $\{1,\ldots,\lfloor p/5\rfloor=26\}$. We are concerned with inference for the means $\mu_{Y,\hat{\jmath}}$ for $\hat{\jmath}$ in $\hat{J}_{R}$. In the figure below, we plot lower and upper confidence bounds for the mobility gains in selected Seattle tracts. We also apply the same exercise to Cleveland to demonstrate heterogeneity in findings between urban areas.
@@
In urban Seattle, we fail to reject the possibility of null effects at 92% of selected tracts. In urban Cleveland, we fail to reject the possibility of null effects in a comparatively small 35% of selected tracts.
@@
$\bar{\mu}_{Y,\hat{J}_{R}}:=\frac{1}{\lfloor p/5\rfloor}\sum_{\hat{\jmath}\in\hat{J}_{R}}\mu_{Y,\hat{\jmath}}$
Their confidence region for $\bar{\mu}_{Y,\hat{J}_{R}}$ lies above zero, implying that they can reject the possibility of a null effect on economic mobility on aggregate. Our analysis differs from that of Andrews et al. (2023) in that we are concerned with constructing a confidence region satisfying simultaneous coverage of the $\mu_{Y,\hat{\jmath}}$. It thereby provides insight into which of the selected tracts most credibly drive the positive aggregate effects.
@@
Figure 3: Proportion of selected tracts for which we fail to reject the possibility of a null effect on economic mobility relative to the average housing voucher recipient, by commuting zone. We provide results for all top-50 CZs by population.
In our final exercise on tract-level effects, we study pairwise comparisons of low and high-opportunity tracts. In particular, we consider the thought experiment of a household moving from an arbitrary low-opportunity to an arbitrary high-opportunity tract, and seek to study how often we can reject the possibility of a null effect on economic mobility associated with this move. As discussed in remark 3.5, this problem is closely related to the problem of inference for ranks. In their application to the economic mobility literature, Mogstad et al. (2023) consider the problem of inference on all pairwise comparisons in order to be agnostic on movement patterns among CMTO enrollees. Our analysis is notably different from that of Mogstad et al. (2023), since we restrict attention to pairwise comparisons of selected tracts. This allows us to be agnostic on movement patterns only among CMTO enrollees living in low-ranked tracts prior to treatment.
@@
CZ $\tau$-best $\tau$-worst % Fail to Reject Lowest LCB Highest LCB TS
two-step two-step two-step
Seattle 26 26 87% [-0.33, 0.59] [0.077, 0.50]
26 1 100% [-0.19, 0.60] [-0.055, 0.66]
1 1 100% [-0.055, 0.66] [-0.055, 0.66]
Cleveland 34 34 31% [-0.076, 0.59] [0.11, 0.49]
34 1 9% [-0.056, 0.38] [0.11, 0.49]
1 1 0% [0.12, 0.48] [0.12, 0.48]
@@
In the table below, we replicate our analysis of the mobility effects of moving from low to high opportunity areas at the commuting zone level. Our results on mobility effects at the commuting zone level are more conclusive than our findings at the tract-zone level. This is in part due to the simple fact that the standard errors on CZ level effects are smaller than those on tract level effects. We find that, for the majority of high and low-opportunity commuting zone pairs, we can indeed conclude that there exists a non-zero difference in the true mobilities of both CZs.
@@
Top Bottom % Fail to Reject Lowest LCB Highest LCB
two-step two-step two-step
50% 50% 18.5% [-0.28, 0.35] [0.33, 0.44]
33% 33% 8.2% [-0.25, 0.50] [0.33, 0.44]
20% 20% 4.7% [-0.19, 0.52] [0.34, 0.44]
10% 10% 3.5% [-0.14, 0.55] [0.34, 0.44]
33% 50% 12.0% [0.28, 0.35] [0.33, 0.44]
20% 50% 8.3% [-0.22, 0.38] [0.33, 0.44]
10% 50% 7.1% [-0.18, 0.41] [0.34, 0.44]
33% 67% 19.9% [-0.28, 0.35] [0.33, 0.44]
20% 80% 19.1% [-0.25, 0.29] [0.33, 0.44]
10% 90% 17.6% [-0.22, 0.31] [0.34, 0.44]
@@
By focusing on selected CZs, however, we are able to distinguish high and low-opportunity commuting zones with high frequency.
@@ 7 Simulation Study
In this section, we conduct an extended simulation study comparing the projection and two-step approaches to inference on multiple winners across a broad range of simulation designs. We demonstrate that the two-step approach to inference on multiple winners outperforms the projection approach almost uniformly. This outperformance is most pronounced in instances where the set of asymptotic winners is a proper subset of the set of all possible selections. In the simulation results presented in this section, we demonstrate that our methods perform favorably relative to standard approaches to inference such as projection, as well as relative to the state of the art of Zrnic and Fithian (2024b) and Zrnic and Fithian (2024a).
@@
In all of our simulations, we consider four distinct covariance cases. In particular, we have a simple covariance case where $X$ and $Y$ are perfectly correlated but the $X_{j}$ are independent, a low covariance case where all units are weakly correlated, a medium covariance case, and a high covariance case. In these four cases, we denote the variance covariance matrices by $\Sigma_{simple},\Sigma_{low},\Sigma_{medium}$, or $\Sigma_{high}$. We provide explicit formulae for these variance covariance matrices in appendix 11. In the simulation results presented in this section, we take $\Sigma=\Sigma_{simple}$. We scale $\Sigma$ by $1/n$ for sample size $n$ equal to 100, 1000,and 10000100,\ 1000,\ $\text${and }10000. We emphasize that in all simulations, we treat $\Sigma$ as known for computational simplicity.
We provide a table including results from selected designs, namely the low correlation case of designs A, B, C, and D which demonstrate the two-step approach’s “clear-winner” property (see designs A and B).
@@
100 1000 10000
A Projection 0.991 0.991 0.992
two-step 0.966 0.955 0.955
Zoom 0.977 0.951 0.950
Locally Simultaneous 0.971 0.955 0.955
B Projection 0.993 0.996 0.995
two-step 0.969 0.970 0.960
Zoom 0.985 0.985 0.971
Locally Simultaneous 0.979 0.976 0.962
C Projection 0.976 0.976 0.976
two-step 0.962 0.962 0.962
Zoom 0.977 0.977 0.977
Locally Simultaneous 0.979 0.979 0.979
D Projection 0.991 0.991 0.991
two-step 0.970 0.970 0.969
Zoom 0.979 0.976 0.976
Locally Simultaneous 0.978 0.978 0.979
@@
We find that, between the models described above, the two-step approach substantially outperforms the projection approach when the set of winners $J_{R}(P)$ is clear, and specifically a proper subset of $J$ such that $J\setminus J_{R}(P)$ is large. Moreover, in intermediate cases when the set of winners is moderately clear, as in designs A and B for low $n$, the two-step approach outperforms the approach of Zrnic and Fithian (2024a), which is based on the zoom test. [^7] Our implementation of the zoom test is based on the step-down procedure outlined in section 4 of Zrnic and Fithian (2024a). In general, the two-step approach also outperforms the approach of Zrnic and Fithian (2024b). Quantitatively, we have that in the four simulations presented above, the two-step approach to inference can reduce absolute overcoverage error by up to 88% relative to projection inference, 50% relative to locally simultaneous inference, and 56% relative to the zoom test. Across all simulations, including those in appendix 11, the two-step approach to inference reduces overcoverage error by up to 96% relative to projection inference, up to 71% relative to locally simultaneous inference, and up to 67% relative to the zoom test.
@@
100 1000 10000
A two-step 0.835 0.780 0.780
Zoom 0.872 0.786 0.763
Locally Simultaneous 0.879 0.781 0.780
B two-step 0.824 0.784 0.730
Zoom 0.900 0.842 0.758
Locally Simultaneous 0.888 0.827 0.731
C two-step 0.937 0.937 0.937
Zoom 1.003 1.003 1.003
Locally Simultaneous 1.014 1.014 1.014
D two-step 0.846 0.846 0.846
Zoom 0.914 0.873 0.873
Locally Simultaneous 0.887 0.886 0.886
@@
Moreover, the table above demonstrates that, over a wide range of data generating processes, the two-step approach to inference provides tighter confidence regions than the projection, zoom, and locally simultaneous approaches to inference. Indeed, the interval lengths of the two-step approach to inference may be up to 27% shorter than projection inference. Moreover, two-step inference may be up to 11% shorter than inversions of the zoom test and up to 8% shorter than locally simultaneous inference.
@@
In general, we recommend choosing $\beta=\alpha/10$. As our simulation results demonstrate, the two-step approach to inference performs quite well under such a choice of $\beta$. This choice of $\beta$ is the same as in Andrews et al. (2023) and Romano et al. (2014).